# Alignment Drift in CEFR-prompted LLMs for Interactive Spanish Tutoring

### v1

@@ 1 Introduction
The popularization of large language models (LLMs), particularly through the emergence of user-friendly interfaces such as ChatGPT, has led many stakeholders across society to consider how to use such technology effectively and safely to facilitate access to knowledge and education. Language education has not been immune to this hype, and with seemingly good cause, since LLMs show potential across a range of areas where they might enhance language learning.
@@ 3 Experimental Design
Data generation (Section 3) and analysis (Sections 4 & 5) were carried out in Python (v3.12), with the exception of running linear mixed effects models in R (v4.4.3). All code and the dataset is available on the GitHub repositories:
  • Generation: [REDACTED FOR REVIEW]
  • Dataset & Analysis: [REDACTED FOR REVIEW]
@@ 3.2 Teacher-Student Dialogue Simulation
We simulated a language tutoring scenario by deploying an LLM with separate chat histories as both the "tutor" and "student". As current LLM systems are stateless (Yu et al., 2025), we could instantiate a single LLM object, and then interchange the chat history, having separate chat histories for the student and the tutor (see the graphical overview in Appendix A.2).
@@
Despite being instructed to "keep everything in Spanish" (Figure 1), a number of models generated non-Spanish text. For instance, Gemma and Llama tended to include English content. This happened primarily for the A1 level, where they sometimes provided English translations in parentheses alongside their Spanish sentences. Also, Qwen occasionally switched mid-generation to Mandarin Chinese. To avoid confounding our analysis, we applied a simple language detection algorithm to the tutor LLM’s outputs using the Python library lingua. [^3] https://github.com/pemistahl/lingua-py If English or Mandarin was detected in any sentence, we re-generated the tutor LLM’s response before continuing the dialogue.
@@ 3.3 System Prompts
We created custom system prompts in English for the tutor LLM. These prompts differed only in key, level-specific phrasing. Along with terms such as "beginner," "intermediate," and "advanced," an additional description of a learner’s abilities at the particular level was provided, taken from the CEFR Global scale (see Section 2.2). Figure 1 shows the system prompt for A1 with the level-specific wording highlighted (prompts for B1 and C1 can be viewn in Appendix A.3).
@@ 4.1 Traditional Readability Metrics
Gutiérrez de Polini is a metric specifically created for Spanish (Gutiérrez de Polini, 1972). Unlike the previous two metrics, it does not rely on syllables, but instead considers the number of characters per words and words per sentences (Vásquez-Rodríguez et al., 2022).
@@ 4.2 Structural Complexity
We extract Text Length of each message, operationalised as the token count, as it is included in the definition of the C1 level in the CEFR Global scale (i.e., the student can understand "a wide range of demanding, longer texts" (Council of Europe, 2025b)). A small study on ChatGPT also showed that the model tended to generate longer texts for higher levels of CEFR (Ramadhani et al., 2023). Moreover, in machine classification studies of texts across languages, text length was considered an important predictor of CEFR level (Bestgen, 2020; Yekrangi, 2022).
@@ 5.1 Readability Metrics
The average readability scores over time are shown for all models and CEFR levels in Figure 2. Across LLMs, scores from all three readability metrics decrease as proficiency increases, with A1 having the highest scores (easier to read) and C1 the lowest scores (harder to read). [^8] As expected (Section 4.1), there is a clear resemblance in scores from Fernández Huerta and Szigriszt Pazos, but it is worth noting that the scores are not identical. However, despite starting from different baselines, all curves slowly decrease in readability over time, reducing the differences between CEFR levels as well. A notable exception is Gemma, which has a sudden spike around the last messages in B1 for the Fernández Huerta and Szigriszt Pazos scores. The same behavior is present but less pronounced for the Gutiérrez de Polini scores.
Despite differences in average scores, the confidence intervals reveal some overlap between the levels. These differ across LLMs with a model such as Qwen having a much greater overlap between levels B1 and C1 than Llama. Both these models also begin with generally higher Fernández Huerta and Szigriszt Pazos scores across levels than Gemma and Mistral.
When examining the full distribution of scores as density plots (Figure 3), the overlap between levels across all models is more evident. The distributions also reveal that a small, but not insignificant, portion of Fernández Huerta/Szigriszt Pazos scores reaches around 50 for C1 for Llama and Gemma. This is well below the average scores, and indicates that the LLMs are capable of producing quite complex text, even if they often do not.
Despite the overlapping scores, all mixed effects models revealed that B1 and C1 $(p<0.001)$ had significantly lower readability scores than the baseline A1 ($\beta_{0}$). Across LLMs, the estimates ($\beta$) for Fernández Huerta ranged between -4 and -9 for B1 and -12 and -17 for C1999Given the nature of mixed effects models, no direct conclusion can be drawn about the significance of the difference between levels B1 and C1, as the tests only evaluate the difference relative to the baseline, A1. (See Appendix A.5.1).
@@ 5.2 Structural Features
The same pattern occurs for the MDD scores for models Llama and Qwen, although with closely intersecting curves for C1 and B1. The results are even more muddled for Gemma and Mistral. These results are reflected in the full distributions. Qwen is an outlier when it comes to text length with a much greater uncertainty in average lengths, having a few generations that reach above 2000 tokens as seen on the density plot, which is far above the other LLMs whose highest generations are around 800-1000 tokens.
@@ 6 Discussion
Nevertheless, even when evaluating the readability metrics, it remains debatable whether the differences between levels is large enough to accurately reflect the intended proficiency levels. With average values ranging between 110 and 70 for the Fernández Huerta scores, the readability is equivalent to Spanish school children, even at an average of 70 (see Appendix A.4). While it is unclear how this translates to L2 learners of Spanish, it could suggest that the LLMs have not managed generate text appropriate for the proficiency levels, at least for the C1 level. Refer to the Limitations for other considerations of the metrics.
Finally, the LLMs did not perform equally, which could help inform the choice of a suitable LLM to serve as an language tutor in Spanish, at least for initial development. A model like Llama is relevant to highlight as a well-performing model although its license might be too restrictive for some applications (Meta, 2024).
@@ 7 Conclusion
While we are primarily interested in the application of LLMs in an educational scenario, we want to stress the importance of additional considerations and evaluation before their real-world deployment. Firstly, we recognize that the models may reflect cultural biases that could be inappropriate for the target student population. Therefore, cultural alignment may be necessary before their implementation (Tao et al., 2024; Li et al., 2024).
Moreover, some of the models may not be properly instruction-tuned to align with human principles (e.g., the removal of toxic content). For instance, Mistral, designed for demonstration purposes, lacks "moderation mechanisms" according to the Mistral AI team (Jiang et al., 2024). Such a model would require further development before being suitable for real-world applications.
@@
Despite the target language being Spanish, we defined the system prompt in English. This might explain why the American multilingual models, Gemma and Llama, were prone to producing English content. However, this does not account for why Qwen occasionally generated Mandarin Chinese despite the absence of the Mandarin in the system prompt. This unintended behavior may instead reflect the composition of the training data, with Qwen likely containing more Chinese-language data [^10] Qwen 2.5’s predecessor, Qwen 7B, has a technical memo stating that most of its training data is ”in English and Chinese.” (Qwen Team, 2023) However, Qwen 2.5’s technical report does not explicitly mention this, aside from emphasizing the model’s evaluation in these two languages (Qwen Team et al., 2025). than the American models, where English likely dominates.
@@ Appendix A Appendix
A.1 Techhnical Details about Inference
All LLM inference was run using the Hugging Face transformers package (Wolf et al., 2020) on a university-developed cloud computing platform. Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, and Qwen-2.5-7B Instruct were run on a single NVIDIA L40 GPU (48 GB), with 96 GB of system memory and 8 vCPUs. Gemma-3-12B-IT was run on a system with double the resources, utilizing two NVIDIA L40 GPUs. Due to the higher resource demands of Gemma-3-12B-IT, we chose to run it with a lower precision (bfloat16). This minor difference in precision from the other models was not considered impactful for the model comparisons.
@@ A.2 Illustration of Simulation Framework
Graphical overview of the simulation framework. The actual simulations consisted of nine rounds, not two. The example text (abbreviated) is taken from a simulated conversation by Mistral in A1. For the implementation in code, refer to the file on GitHub: [REDACTED FOR REVIEW].
@@ A.4 Interpretation of Readability Scales
Due to slight differences in reporting, we provide two variations of the interpretation tables for Fernández Huerta and Szigriszt Pazos. While such tables are commonly reported for the two metrics, interpretations of Gutiérrez de Polini were difficult to find beyond the version proposed by Scott (2024b).
Fernández Huerta Szigriszt Pazos Level
@@
Szigriszt Pazos Level Spanish Grade Level US Grade Level Age Group

### v2

@@ 1 Introduction
The popularization of large language models (LLMs), particularly through the emergence of user-friendly interfaces such as ChatGPT, has led many stakeholders across society to consider how to use such technology effectively and safely to facilitate access to knowledge and education (Yan et al., 2024). Language education has not been immune to this hype, and with seemingly good cause, since LLMs show potential across a range of areas where they might enhance language learning.
@@ 3 Experimental Design
Data generation (Section 3) and analysis (Sections 4 & 5) were carried out in Python (v3.12.3), with the exception of running linear mixed effects models in R (v4.4.3). All code and the dataset is available on the GitHub repositories:
@@ 3.2 Teacher-Student Dialogue Simulation
We simulated a language tutoring scenario by deploying an LLM with separate chat histories as both the "tutor" and "student". Current LLM systems are stateless (Yu et al., 2025), with the entire chat history being processed by the model during each interaction. This allowed us to instantiate a single LLM object, and then interchange the chat history, maintaining one history for the student and another for the tutor (see the graphical overview in Appendix A.2).
@@
Despite being instructed to "keep everything in Spanish" (Figure 1), a number of models generated non-Spanish text. [^3] We also discuss this in a subsection of the Limitations. For instance, Gemma and Llama tended to include English content. This happened primarily for the A1 level, where they sometimes provided English translations in parentheses alongside their Spanish sentences. Also, Qwen occasionally switched mid-generation to Mandarin Chinese. To avoid confounding our analysis, we applied a simple language detection algorithm to the tutor LLM’s outputs using the Python library lingua. [^4] https://github.com/pemistahl/lingua-py If English or Mandarin was detected in any sentence, we re-generated the tutor LLM’s response before continuing the dialogue.
@@ 3.3 System Prompts
We created custom system prompts in English for the tutor LLM. These prompts differed only in key, level-specific phrasing. Along with terms such as "beginner," "intermediate," and "advanced," an additional description of a learner’s abilities at the particular level was provided, taken from the CEFR Global scale (see Section 2.2). Figure 1 shows the system prompt for A1 with the level-specific wording highlighted (prompts for B1 and C1 can be viewed in Appendix A.3).
@@ 4.1 Traditional Readability Metrics
Gutiérrez de Polini is a metric specifically created for Spanish (Gutiérrez de Polini, 1972). Unlike the previous two metrics, it does not rely on syllables, but instead considers the number of characters per word and words per sentence (Vásquez-Rodríguez et al., 2022).
@@ 4.2 Structural Complexity
We extract Text Length of each message, operationalized as the token count, as it is included in the definition of the C1 level in the CEFR Global scale (i.e., the student can understand "a wide range of demanding, longer texts" (Council of Europe, 2025b)). A small study on ChatGPT also showed that the model tended to generate longer texts for higher levels of CEFR (Ramadhani et al., 2023). Moreover, in machine classification studies of texts across languages, text length was considered an important predictor of CEFR level (Bestgen, 2020; Yekrangi, 2022).
@@ 5.1 Readability Metrics
The average readability scores over time are shown for all models and CEFR levels in Figure 2. Across LLMs, scores from all three readability metrics decrease as proficiency increases, with A1 having the highest scores (easier to read) and C1 the lowest scores (harder to read). [^9] As expected (Section 4.1), there is a clear resemblance in scores from Fernández Huerta and Szigriszt-Pazos, but it is worth noting that the scores are not identical. However, despite starting from different baselines, all curves slowly decrease in readability over time, reducing the differences between CEFR levels as well. A notable exception is Gemma, which has a sudden spike around the last messages in B1 for the Fernández Huerta and Szigriszt-Pazos scores. The same behavior is present but less pronounced for the Gutiérrez de Polini scores.
Despite differences in average scores, the confidence intervals reveal some overlap between the levels. These differ across LLMs with a model such as Qwen having a much greater overlap between levels B1 and C1 than Llama. Both these models also begin with generally higher Fernández Huerta and Szigriszt-Pazos scores across levels than Gemma and Mistral.
When examining the full distribution of scores as density plots (Figure 3), the overlap between levels across all models is more evident. The distributions also reveal that a small, but not insignificant, portion of Fernández Huerta/Szigriszt-Pazos scores reaches around 50 for C1 for Llama and Gemma. This is well below the average scores, and indicates that the LLMs are capable of producing quite complex text, even if they often do not.
Despite the overlapping scores, all mixed effects models revealed that B1 and C1 $(p<0.001)$ had significantly lower readability scores than the baseline A1 ($\beta_{0}$). Across LLMs, the estimates ($\beta$) for Fernández Huerta ranged between -4 and -9 for B1 and -12 and -17 for C1101010Given the nature of mixed effects models, no direct conclusion can be drawn about the significance of the difference between levels B1 and C1, as the tests only evaluate the difference relative to the baseline, A1. (See Appendix A.5.1).
@@ 5.2 Structural Features
The same pattern occurs for the MDD scores for Llama and Qwen, although with closely intersecting curves for C1 and B1. The results are even more muddled for Gemma and Mistral. These results are reflected in the full distributions. Qwen is an outlier when it comes to text length with a much greater uncertainty in average lengths, having a few generations that reach above 2000 tokens as seen on the density plot, which is far above the other LLMs whose highest generations are around 800-1000 tokens.
@@ 6 Discussion
Nevertheless, even when evaluating the readability metrics, it remains debatable whether the differences between levels are large enough to accurately reflect the intended proficiency levels. With average values ranging between 110 and 70 for the Fernández Huerta scores, the readability is equivalent to Spanish school children, even at an average of 70 (see Appendix A.4). While it is unclear how this translates to L2 learners of Spanish, it could suggest that the LLMs have not managed generate text appropriate for the proficiency levels, at least for the C1 level. Refer to the Limitations for other considerations of the metrics.
An additional concern is that the observed alignment drift could have been driven by a possible drift in the student LLM (i.e., the tutor adapting to the student and vice versa). As we neither optimized nor examined the student LLM, it remains unclear how this influenced the outcome or how this would differ with human users. However, LLMs have also shown difficulty in following system prompts over the course of multi-turn dialogue in other domains with real user messages (Qiu and Yang, 2024). Hence, we do not expect a substantial difference between using human or LLM students given our current framework. We leave it to future research to investigate the exact influence of the student LLM on the tutor LLM’s alignment drift, potentially including human students as a point of comparison.
As a final remark, we note that the LLMs did not perform equally, which could help inform the choice of a suitable LLM to serve as an language tutor in Spanish, at least for initial development. A model like Llama is relevant to highlight as a well-performing model although its license might be too restrictive for some applications (Meta, 2024).
@@ 7 Conclusion
We wish to stress the importance of additional considerations and evaluation of LLMs before their real-world deployment in educational contexts. Firstly, we recognize that the models may reflect cultural biases that could be inappropriate for the target student population. Therefore, cultural alignment may be necessary before their implementation (Tao et al., 2024; Li et al., 2024). Moreover, some of the models may not be properly instruction-tuned to align with human principles (e.g., the removal of toxic content). For instance, Mistral, designed for demonstration purposes, lacks "moderation mechanisms" according to the Mistral AI team (Jiang et al., 2024). Such a model would require further development before being suitable for real-world applications.
These ethical concerns are increasingly urgent when considering the impact that generative AI may have on language learners. For example, L2 learners might over-rely on ChatGPT (Yang and Li, 2024) such as using it to write complete assignments rather than as a supplementary tool (Yan, 2023). More broadly, the ELIZA effect (Weizenbaum, 1966), describing our tendency to attribute human-like qualities such as "understanding" to machines (Mitchell and Krakauer, 2023), may contribute problematically to the overtrust of AI chatbots (Reinecke et al., 2025). We urge developers to prioritize the responsible implementation of LLM systems for education and believe that our research contributes to work in this direction.
@@
Despite the target language being Spanish, we defined the system prompt in English. This might explain why the American multilingual models, Gemma and Llama, were prone to producing English content. However, this does not account for why Qwen occasionally generated Mandarin Chinese despite the absence of Mandarin in the system prompt. This unintended behavior may instead reflect the composition of the training data, with Qwen likely containing more Chinese-language data [^11] Qwen 2.5’s predecessor, Qwen 7B, has a technical memo stating that most of its training data is ”in English and Chinese.” (Qwen Team, 2023). However, Qwen 2.5’s technical report does not explicitly mention this, aside from including evaluation on these two languages (Qwen Team et al., 2025). than the American models, where English likely dominates.
@@
Acknowledgments
All of the computation done for this project was performed on the UCloud interactive HPC system, which is managed by the eScience Center at the University of Southern Denmark.
@@ Appendix A Appendix
A.1 Technical Details about Inference
All LLM inference was run using the Hugging Face transformers package (Wolf et al., 2020) on a cloud-based interactive HPC platform (Python v3.12.3, Ubuntu v24.04). Llama, Mistral, and Qwen were run on a single NVIDIA L40 GPU (48 GB), with 96 GB of system memory and 8 vCPUs, while Gemma was run on a system utilizing two NVIDIA L40 GPUs. Due to the higher resource demands of Gemma, we chose to run it with a lower precision (bfloat16). This minor difference in precision from the other models was not considered impactful for the model comparisons.
@@ A.2 Illustration of Simulation Framework
Graphical overview of the simulation framework. The actual simulations consisted of nine rounds, not two. The example text (abbreviated) is taken from a simulated conversation by Mistral in A1. For the implementation in code, refer to the file on GitHub: INTERACT-LLM/Interact-LLM/src/scripts/alignment_drift/simulate.py.
@@ A.4 Interpretation of Readability Scales
Due to slight differences in reporting, we provide two variations of the interpretation tables for Fernández Huerta and Szigriszt-Pazos. While such tables are commonly reported for the two metrics, interpretations of Gutiérrez de Polini were difficult to find beyond the version proposed by Scott (2024b).
Fernández Huerta Szigriszt-Pazos Level
@@
Szigriszt-Pazos Level Spanish Grade Level US Grade Level Age Group